whyjustwhy:

can some1 help me 1+4-(-48)= 23

xcoledd1:

umm

xcoledd1:

how can i help

Ultrilliam:

What is the question supposed to be?

Midnight97:

Let me see what I can do

xcoledd1:

@whyjustwhy

xcoledd1:

it is not true if ur trying to ask

xcoledd1:

it is equal to umm 53

Midnight97:

A negative and a negative make a positive. So you can rewrite the problem 1+4+48=23 then you need to add 1,4 and 48 that'll be 58=23 that'll be your answer

Midnight97:

Hope it helps

Ultrilliam:

@midnight97 wrote:
A negative and a negative make a positive. So you can rewrite the problem 1+4+48=23 then you need to add 1,4 and 48 that'll be 58=23 that'll be your answer
That isn't true in this case since we aren't multiplying, normally we would multiply so I can understand the confusion, but in this instance the parenthesis are actually pointless, and probably only there to help visualize that it is subtracting -48.

Midnight97:

@ultrilliam wrote:
@midnight97 wrote:
A negative and a negative make a positive. So you can rewrite the problem 1+4+48=23 then you need to add 1,4 and 48 that'll be 58=23 that'll be your answer
That isn't true in this case since we aren't multiplying, normally we would multiply so I can understand the confusion, but in this instance the parenthesis are actually pointless, and probably only there to help visualize that it is subtracting -48.
Okay well I tried

Ultrilliam:

You are partially correct though, the steps to solving it are pretty much 1+4+48. But your answer is incorrect.

Midnight97:

I understand

whyjustwhy:

i thought it was a fake math problem -_-

HotPockets14:

It falso the answer to that is not 23